### v1

# Universal bound states and resonances with Coulomb plus short-range potentials

Abstract
We study charged particles in three dimensions interacting via a short-range potential in addition to the Coulomb potential. When the Bohr radius and the scattering length are much larger than the potential range, low-energy physics of the system becomes independent from details of the short-range potential. We develop the zero-range theory to describe such universal physics in terms of the Bohr radius and the scattering length by generalizing the Bethe-Peierls boundary condition based on the self-adjoint extension of the Hamiltonian, which is then applied to reveal bound states and resonances of two charged particles resulting as poles in the complex energy plane of the $S$ matrix. In particular, infinite resonances are found for a repulsive Coulomb potential, one of which turns into a bound state with increasing inverse scattering length, whereas infinite bound states exist for an attractive Coulomb potential with no resonances at any scattering length.
I Introduction
When particles interact via a short-range potential with its scattering length much larger than the potential range, low-energy physics of the system becomes universal, i.e., independent from details of the short-range potential [1]. Such universal physics is parametrized by the scattering length only and can be relevant to diverse systems in physics due to the universality. One representative example is the BCS-BEC crossover for many fermions [2, 3, 4, 5], which has attracted considerable interests not only from ultracold atoms but also from nuclear physics [6, 7]. Another representative example is the Efimov effect for few bosons [8, 9, 1, 10, 11, 12, 13, 14, 15], which was theoretically predicted in nuclear physics [16, 17], experimentally observed in ultracold atoms [18], and potentially expected in condensed matter systems [19, 20, 21].
A theoretical framework to describe the universal physics in terms of the scattering length is called the zero-range theory, which can be implemented in various ways [1]. One naive way is to take the zero-range limit of some finite-range potential under fixed scattering length, whereas a zero-range potential can directly be constructed with the Huang-Yang pseudopotential being the most famous among other equivalent representations [22, 23]. Instead of providing the Hamiltonian with potentials, the zero-range theory can also be implemented by imposing the so-called Bethe-Peierls boundary condition on the wave function,
$\lim_{r\to 0}\psi(r)\propto\frac{1}{r}-\frac{1}{a}+O(r),$ (1)
allowing the singularity determined by the scattering length $a$ [24].
The above zero-range theory assumes a short-range potential only, so that it cannot be applied to charged particles interacting via the Coulomb potential in addition to a short-range potential. Such charged particles and resulting bound states and resonances, however, have been important in diverse fields such as atomic, molecular, and chemical physics [25, 26, 27, 28], nuclear and hadron physics [29, 30, 31, 32, 33], and dark-matter astrophysics [34]. Therefore, it is worthwhile to develop the zero-range theory in the presence of a Coulomb potential and reveal universal aspects of two charged particles interacting via a short-range potential, which will be presented in Secs. II and III of this paper, respectively.
II Zero-range theory
II.1 Self-adjoint extension
Let us study two charged particles in three dimensions, whose relative wave function in the $s$-wave channel obey $\hat{H}r\psi(r)=Er\psi(r)$ with
$\hat{H}=-\frac{\hbar^{2}}{2m}\frac{d^{2}}{dr^{2}}\pm\frac{\hbar^{2}}{ma_{0}r}.$ (2)
Here, $r\in(0,\infty)$ is the interparticle separation, $m$ is the reduced mass, $a_{0}$ is the Bohr radius, and the upper (lower) sign corresponds to the repulsive (attractive) Coulomb potential throughout this section. We wish to implement the zero-range interaction by finding appropriate boundary conditions on the wave function at $r\to 0$, which can be achieved based on the self-adjoint extension of the Hamiltonian. Because its detailed accounts can be found for example in Refs. [35, 36, 37], only the essentials are described below.
We introduce $\chi(r)=\sqrt{4\pi}r\psi(r)$ assumed to be square integrable on $(0,\infty)$. The Hamiltonian has to be hermitian, meaning that the right-hand side of
[…]
$=\frac{\hbar^{2}}{2m}\lim_{r\to 0}W[\varphi^{*}(r),\chi(r)]$ (3)
vanishes with $W[\varphi^{*}(r),\chi(r)]=\varphi(r)^{*}\chi^{\prime}(r)-\varphi^{\prime}(r)^{*}\chi(r)$ being the Wronskian. Furthermore, the Hamiltonian has to be self-adjoint, meaning that the maximal domain of $\varphi$ on which $\hat{H}$ can act coincides with the given domain of $\chi$. For example, if $\lim_{r\to 0}\chi(r)=\lim_{r\to 0}\chi^{\prime}(r)=0$ are imposed, $\hat{H}$ is hermitian without any conditions on $\varphi$, violating its self-adjointness. On the other hand, if only $\lim_{r\to 0}\chi(r)=0$ is imposed, $\lim_{r\to 0}\varphi(r)=0$ is required in order for $\hat{H}$ to be hermitian, making it self-adjoint as well. The latter case is the usual boundary condition with no zero-range interaction implemented.
The self-adjoint Hamiltonian actually admits a broader class of boundary conditions at $r\to 0$. To see this, we express the Wronskian in Eq. (II.1) as
$W[\varphi^{*}(r),\chi(r)]=A[\varphi(r)]^{*}B[\chi(r)]-B[\varphi(r)]^{*}A[\chi(r)],$ (4)
where
$A[\chi(r)]$ $\equiv W[\chi(r),f(r)\cos\delta-g(r)\sin\delta],$ (5)
$B[\chi(r)]$ $\equiv W[\chi(r),f(r)\sin\delta+g(r)\cos\delta]$ (6)
are defined with $f(r)$ and $g(r)$ being arbitrary real functions subject to $W[f(r),g(r)]=1$ and $\delta$ an arbitrary real constant [38]. Therefore, if $\lim_{r\to 0}A[\chi(r)]=\lim_{r\to 0}A[\varphi(r)]=0$ are imposed, the Hamiltonian is hermitian and self-adjoint. Although $\lim_{r\to 0}B[\chi(r)]=\lim_{r\to 0}B[\varphi(r)]=0$ are also possible, they are redundant because of $B[\chi(r)]=A[\chi(r)]|_{\delta\to\delta-\pi/2}$.
In order for the boundary condition to allow nonvanishing solutions, it is appropriate to choose the auxiliary functions as two independent solutions to $\hat{H}f(r)=\hat{H}g(r)=0$ [38]. We then find
$f(r)$ $=\sqrt{r}\,I_{1}\!\left(2\sqrt{\frac{2r}{a_{0}}}\right),$ (7)
$g(r)$ $=-2\sqrt{r}\,K_{1}\!\left(2\sqrt{\frac{2r}{a_{0}}}\right)$ (8)
for the repulsive Coulomb potential and
$f(r)$ $=\sqrt{r}\,J_{1}\!\left(2\sqrt{\frac{2r}{a_{0}}}\right),$ (9)
$g(r)$ $=\pi\sqrt{r}\,Y_{1}\!\left(2\sqrt{\frac{2r}{a_{0}}}\right)$ (10)
for the attractive Coulomb potential in terms of the regular and singular (modified) Bessel functions [39]. Because $\lim_{r\to 0}A[\chi(r)]=0$ implies $\chi(r)$ proportional to their specific superposition of $f(r)\cot\delta-g(r)$ at $r\to 0$, the boundary condition reads
$\lim_{r\to 0}\chi(r)\propto 1+\left[\cot\delta\pm\ln\!\left(e^{2\gamma-1}\frac{2r}{a_{0}}\right)\right]\frac{2r}{a_{0}}+O(r^{2}\ln r),$ (11a)
which constitutes the generalization of the Bethe-Peierls boundary condition in the presence of a Coulomb potential. The resulting boundary condition is parametrized by $\cot\delta\in(-\infty,\infty)$, which in comparison to Eq. (1) may be denoted as
$\cot\delta=-\frac{a_{0}}{2\tilde{a}}.$ (11b)
Here, $\tilde{a}$ is the analog of the scattering length in the presence of a Coulomb potential, and Eq. (1) is indeed recovered from Eq. (11) in the limit of $a_{0}\to\infty$ where the Coulomb potential disappears and $\tilde{a}\to a$.
II.2 Relation to the effective-range expansion
It is instructive to clarify our zero-range interaction in relation to the effective-range expansion in the presence of a Coulomb potential. When a short-range potential exists only at $r\leq R$, the outer wave function solving the radial Schrödinger equation (2) for $E=\hbar^{2}k^{2}/(2m)$ is generally provided by
$\chi(r)|_{r>R}=C_{\eta}[F_{\eta}(\rho)\cot\tilde{\delta}(k)+G_{\eta}(\rho)].$ (12)
Here, $F_{\eta}(\rho)$ and $G_{\eta}(\rho)$ are the regular and singular Coulomb wave functions in the $s$-wave channel with $\rho=kr$, $\eta=\pm 1/(ka_{0})$, and $C_{\eta}=\sqrt{2\pi\eta/(e^{2\pi\eta}-1)}$ [39], whereas $\tilde{\delta}(k)$ is the phase shift dependent on the short-range potential. The effective-range expansion then reads
$C_{\eta}^{2}k\cot\tilde{\delta}(k)+2k\eta h_{\eta}=-\frac{1}{\tilde{a}}+\frac{\tilde{r}}{2}k^{2}+O(k^{4}),$ (13)
where $h_{\eta}=[\Psi(i\eta)+\Psi(-i\eta)]/2+\ln(ka_{0})$ with $\Psi(z)$ being the digamma function and $\tilde{r}$ is the analog of the effective range in the presence of a Coulomb potential [29, 30, 31] (see also Ref. [33] and references therein).
Since $R=0$ for the zero-range interaction (ZRI), Eq. (12) extends down to $r\to 0$, where the boundary condition in Eq. (11) determines the phase shift as a function of $k$. With the power-series expansions of the Coulomb wave functions in small $\rho$, [^1] Specifically, $F_{\eta}(\rho)=C_{\eta}\rho+O(\rho^{2})$ and $C_{\eta}G_{\eta}(\rho)=1+\eta\left[\Psi(i\eta)+\Psi(-i\eta)+4\gamma-2+2\ln(2\rho)\right]\rho+O(\rho^{2}\ln\rho)$ for $\rho\to 0$ are employed. we find
$C_{\eta}^{2}k\cot\tilde{\delta}(k)+2k\eta h_{\eta}\Big{|}_{\text{ZRI}}=-\frac{1}{\tilde{a}},$ (14)
so that the effective range and all the higher-order coefficients vanish. Therefore, the zero-range interaction is parametrized by the scattering length only as expected, which is consistent with the causality bound requiring $\tilde{r}\leq 0$ for $R=0$ [33].
II.3 Zero-range limit of a finite-range potential
Although the boundary condition in Eq. (11) is suitable for practical purposes, the zero-range interaction can also be implemented by taking the zero-range limit of a finite-range potential under fixed scattering length. For demonstration, we employ a toy potential,
$V(r)=\begin{cases}\displaystyle-\frac{\hbar^{2}v^{2}}{2m}&(r\leq R)\\[8.0pt] \displaystyle\pm\frac{\hbar^{2}}{ma_{0}r}&(r>R),\end{cases}$ (15)
where the Coulomb potential is replaced by a square-well potential at $r\leq R$. The resulting scattering length can be determined by connecting the zero-energy solution $\chi(r)|_{r\leq R}\propto\sin(vr)$ to $\chi(r)|_{r>R}\propto f(r)\cot\delta-g(r)$ at $r=R$, leading to
$v\cot(vR)=\frac{f^{\prime}(R)\cot\delta-g^{\prime}(R)}{f(R)\cot\delta-g(R)}$ (16)
with $f(r)$ and $g(r)$ provided by Eqs. (7)–(10) and $\cot\delta$ by Eq. (11b). The zero-range interaction is achieved by taking the limit of $R\to 0$, where $\tilde{a}$ is fixed by tuning the potential depth according to
$v=\frac{\pi}{2R}+\frac{2}{\pi\tilde{a}}\pm\frac{4}{\pi a_{0}}\ln\!\left(e^{-2\gamma}\frac{a_{0}}{2R}\right)+O(R\ln^{2}\!R)$ (17)
at its least value. Therefore, the zero-range interaction is always attractive, whose attraction becomes stronger (weaker) with increasing (decreasing) inverse scattering length, and $\tilde{a}$ depends not only on the short-range potential but also on the Coulomb potential. In particular, $v=\pi/(2R)+2/(\pi a)+O(R)$ is recovered in the limit of $a_{0}\to\infty$, so that $1/\tilde{a}=1/a\mp(2/a_{0})\ln[e^{-2\gamma}a_{0}/(2R)]+O(R\ln^{2}\!R)$ holds for the square-well potential.
III Bound states and resonances
III.1 Scattering matrix
Let us apply the zero-range theory developed in the previous section to reveal bound states and resonances of two charged particles resulting as poles in the complex energy plane of the $S$ matrix. To this end, we start with the wave function solving the radial Schrödinger equation (2) for $E=\hbar^{2}k^{2}/(2m)$,
$\chi(r)=C_{\eta}[F_{\eta}(\rho)\cot\tilde{\delta}(k)+G_{\eta}(\rho)],$ (18)
with the boundary condition in Eq. (11) leading to the phase shift already determined as Eq. (14). Its asymptotic form at large $\rho$ is found to be
$\lim_{r\to\infty}\chi(r)$ $\propto e^{-i[\rho-\eta\ln(2\rho)]}$
$\quad-e^{2i[\sigma_{\eta}+\tilde{\delta}(k)]}e^{i[\rho-\eta\ln(2\rho)]}+O(\rho^{-1}),$ (19)
where $\sigma_{\eta}=[\ln\Gamma(1+i\eta)-\ln\Gamma(1-i\eta)]/(2i)$ is the Coulomb phase shift. [^2] Here, $F_{\eta}(\rho)=\sin[\rho-\eta\ln(2\rho)+\sigma_{\eta}]+O(\rho^{-1})$ and $G_{\eta}(\rho)=\cos[\rho-\eta\ln(2\rho)+\sigma_{\eta}]+O(\rho^{-1})$ for $\rho\to\infty$ are employed [39]. The coefficient of the outgoing wave with respect to the incoming wave defines the $S$ matrix $S(k)=e^{2i[\sigma_{\eta}+\tilde{\delta}(k)]}$ [40], which can be expressed as
$S(k)=\frac{\Gamma(1+i\eta)}{\Gamma(1-i\eta)}\frac{-\frac{1}{\tilde{a}}-2k\eta h_{\eta}+C_{\eta}^{2}ik}{-\frac{1}{\tilde{a}}-2k\eta h_{\eta}-C_{\eta}^{2}ik}$ (20)
in terms of the Bohr radius and the scattering length. Furthermore, the identifies  [^3] They can be obtained with $\Psi(z)=\Psi(1+z)-1/z$ and $\Psi(z)=\Psi(1-z)-\pi\cot(\pi z)$ [39].
$h_{\eta}$ $=\Psi(1-i\eta)+\ln(+ika_{0})+\frac{1-C_{\eta}^{2}}{2i\eta}$ (21)
$=\Psi(1+i\eta)+\ln(-ika_{0})-\frac{1-C_{\eta}^{2}}{2i\eta}$ (22)
lead to
$S(k)=\frac{\Gamma(1+i\eta)}{\Gamma(1-i\eta)}\frac{-\frac{1}{\tilde{a}}-2k\eta[\Psi(1-i\eta)+\ln(+ika_{0})]+ik}{-\frac{1}{\tilde{a}}-2k\eta[\Psi(1+i\eta)+\ln(-ika_{0})]-ik}$ (23)
with $\eta=\pm 1/(ka_{0})$ for repulsive (upper sign) and attractive (lower sign) Coulomb potentials.
Whereas $|S(k)|=1$ holds for real $k$, the $S$ matrix analytically continued to complex $k$ may have poles [32, 40]. In particular, poles at $\operatorname{Re}k=0$ with $\operatorname{Im}k>0$ ($\operatorname{Im}k<0$) correspond to bound (virtual) states whose wave functions decay (grow) exponentially at $r\to\infty$. On the other hand, poles at $\operatorname{Im}k<0$ with $\operatorname{Re}k>0$ ($\operatorname{Re}k<0$) are called resonances (antiresonances) with complex energies, which always appear in pairs because of $S(k)^{*}=S(-k^{*})$.
We note that $\Gamma(1+i\eta)$ has poles at $i\eta=-n$ with $n=1,2,3,\dots$ corresponding to infinite virtual (bound) states for a repulsive (attractive) Coulomb potential with no short-range potential. They are not, however, poles of the $S$ matrix in Eq. (23) because $\lim_{i\eta\to-n}\Gamma(1+i\eta)/\Psi(1+i\eta)=(-1)^{n}/\Gamma(n)$ is finite. Similarly, $\Psi(1-i\eta)$ does not produce poles of the $S$ matrix. Therefore, the $S$ matrix has poles only at $k$ satisfying
$ik+\frac{1}{\tilde{a}}+2k\eta[\Psi(1+i\eta)+\ln(-ika_{0})]=0,$ (24)
with $k=i/a$ recovered in the limit of $a_{0}\to\infty$ where the Coulomb potential disappears and $\tilde{a}\to a$. Its solutions in the complex $k$ plane are now analyzed as functions of $a_{0}/\tilde{a}\in(-\infty,\infty)$ for repulsive and attractive Coulomb potentials separately. In particular, $\arg k\in(-\pi/2,3\pi/2)$ is considered because virtual states are not allowed by Eq. (24).
Figure 1: Complex energies $E=\hbar^{2}k^{2}/(2m)$ of the lowest four and extra resonances in the complex plane of $ka_{0}$ as functions of the inverse scattering length $a_{0}/\tilde{a}$ normalized by the Bohr radius. Several values of $a_{0}/\tilde{a}$ are indicated on their trajectories. The corresponding antiresonances exist on the side of $\operatorname{Re}k<0$ (not shown).
III.2 Repulsive Coulomb potential
For a repulsive Coulomb potential, Eq. (24) with $\eta=+1/(ka_{0})$ reads
$ika_{0}+\frac{a_{0}}{\tilde{a}}+2\left[\Psi\!\left(1+\frac{i}{ka_{0}}\right)+\ln(-ika_{0})\right]=0,$ (25)
which is found to support infinite resonances and corresponding antiresonances at an arbitrary scattering length. Their complex energies in the form of complex $k$ are shown in Fig. 1 as functions of $a_{0}/\tilde{a}$, where two distinct classes of solutions can be seen. One class consists of infinite resonances with $n=1,2,3,\dots$, which appear out of the virtual Rydberg levels and eventually converge into the same levels as
$ka_{0}=-\frac{i}{n}-\frac{2i\tilde{a}}{n^{2}(a_{0}-2\pi i\tilde{a})}+O(\tilde{a}^{2})$ (26)
in the limits of $a_{0}/\tilde{a}\to\pm\infty$. The other class consists of only one resonance, which appears out of
$ka_{0}=\frac{ia_{0}}{\tilde{a}}+2i\ln\left(-e^{-\gamma}\frac{a_{0}}{\tilde{a}}\right)+2\pi+O(\tilde{a}\ln\tilde{a})$ (27)
in the limit of $a_{0}/\tilde{a}\to-\infty$ and then turns into a bound state right at $a_{0}/\tilde{a}\to 0$,
$ka_{0}=\sqrt{-\frac{6a_{0}}{\tilde{a}}}+O(\tilde{a}^{-3/2}),$ (28)
whose binding energy eventually diverges as
$ka_{0}=\frac{ia_{0}}{\tilde{a}}+2i\ln\left(e^{-\gamma}\frac{a_{0}}{\tilde{a}}\right)+O(\tilde{a}\ln\tilde{a})$ (29)
in the limit of $a_{0}/\tilde{a}\to\infty$.
Such a resonance with $\operatorname{Re}k\gg-\operatorname{Im}k$ may be observed as a characteristic structure in the scattering cross section [32, 40], which in the $s$-wave channel is provided by
$\sigma(k)=\frac{\pi}{k^{2}}|S(k)-1|^{2}.$ (30)
The resulting cross sections with respect to that with no short-range potential are shown in Fig. 2 as functions of $ka_{0}$ for several choices of $a_{0}/\tilde{a}=-1,-0.5,-0.1$. With increasing inverse scattering length on its negative side, a sharp rise of the cross section is indeed developed at the resonance energy.
Figure 2: Scattering cross sections in Eq. (30) with $\sigma_{0}=\pi|e^{2i\sigma_{\eta}}-1|^{2}/k^{2}$ subtracted as functions of $k=\sqrt{2mE/\hbar^{2}}$ normalized by the Bohr radius $a_{0}$. The three curves correspond to different inverse scattering lengths $a_{0}/\tilde{a}=-1,-0.5,-0.1$ as indicated by the plot labels.
III.3 Attractive Coulomb potential
For an attractive Coulomb potential, Eq. (24) with $\eta=-1/(ka_{0})$ reads
$ika_{0}+\frac{a_{0}}{\tilde{a}}-2\left[\Psi\!\left(1-\frac{i}{ka_{0}}\right)+\ln(-ika_{0})\right]=0,$ (31)
which is found to support infinite bound states at an arbitrary scattering length. Their binding energies are increasing functions of the inverse scattering length, which in the form of $\kappa=-ik>0$ are shown in Fig. 3 as functions of $a_{0}/\tilde{a}$. Here, all the solutions appear out of the Rydberg levels with $n=1,2,3,\dots$,
$\kappa a_{0}=\frac{1}{n}-\frac{2\tilde{a}}{n^{2}a_{0}}+O(\tilde{a}^{2}),$ (32)
in the limit of $a_{0}/\tilde{a}\to-\infty$. Eventually, the ground-state energy ($n=1$) diverges as
$\kappa a_{0}=\frac{a_{0}}{\tilde{a}}-2\ln\left(e^{-\gamma}\frac{a_{0}}{\tilde{a}}\right)+O(\tilde{a}\ln\tilde{a})$ (33)
in the limit of $a_{0}/\tilde{a}\to\infty$, whereas each excited-state energy ($n=2,3,\dots$) converge into one lower Rydberg level as
$\kappa a_{0}=\frac{1}{n-1}-\frac{2\tilde{a}}{(n-1)^{2}a_{0}}+O(\tilde{a}^{2}).$ (34)
We also note that the binding energies of higher excited states for $n\gg 1$ are well approximated by
$\kappa a_{0}\simeq\frac{1}{n+\frac{1}{\pi}\operatorname{arccot}\!\left(\frac{a_{0}}{2\pi\tilde{a}}\right)}$ (35)
with $\operatorname{arccot}(x)$ defined continuously in the range of $(-\pi,0)$. In contrast to such bound states, no resonances and antiresonances are found at any scattering length.
Figure 3: Binding energies $E=-\hbar^{2}\kappa^{2}/(2m)$ of the lowest five bound states in the form of $\kappa a_{0}$ as functions of the inverse scattering length $a_{0}/\tilde{a}$ normalized by the Bohr radius. The horizontal dotted lines at $\kappa a_{0}=1/n$ indicate the Rydberg levels for $n=1,2,\dots,5$.
IV Summary and outlook
In summary, we studied charged particles in three dimensions interacting via a short-range potential in addition to the Coulomb potential. When the Bohr radius $a_{0}$ and the scattering length $\tilde{a}$ are much larger than the potential range $R$, low-energy (or long-wavelength $2\pi/k$) physics of the system becomes independent from details of the short-range potential. The zero-range theory developed in this paper is suitable to describe such universal physics for $a_{0},\tilde{a},k^{-1}\gg R$ directly in terms of the Bohr radius and the scattering length. Whereas it was applied to two charged particles to reveal infinite resonances (bound states) for a repulsive (attractive) Coulomb potential and their trajectories in the complex energy plane, the zero-range theory can be applied to $N$ charged particles as well. Namely, their wave function is made to obey the $N$-body Schrödinger equation without a short-range potential but with the boundary condition in Eq. (11) imposed whenever two coordinates contact. Such generalization of the Bethe-Peierls boundary condition was obtained based on the self-adjoint extension of the Hamiltonian.
Hopefully, our theoretical framework is to further reveal novel universal physics of few or many particles interacting via Coulomb plus short-range potentials. In particular, an infinite scattering length may be interesting because the interaction scale is provided by the Bohr radius only in spite of the nontrivial short-range potential. Resulting consequences are potentially relevant to diverse systems in atomic, molecular, and chemical physics, nuclear and hadron physics, and dark-matter astrophysics due to the universality.
Acknowledgements.
The authors thank Tetsuo Hyodo, Tomona Kinugawa, and Takuma Yamashita for valuable discussions. This work was supported by JSPS KAKENHI Grant No. JP21K03384.